OpenStudy (anonymous):

medal and fan. what is the equation of the given line in standard form? use integer coefficients. y=-1.7+8.5

OpenStudy (anonymous):

I think you mean y = -1.7x + 8.5?

OpenStudy (anonymous):

yes @nonon

OpenStudy (anonymous):

ok, so standard form is ax + by = c. so manipulate the equation: 1.7x + y = 8.5

OpenStudy (anonymous):

okay, what's our next step @nonon

OpenStudy (anonymous):

Then, multiply all of it by 10 so the coefficients would be integers

OpenStudy (anonymous):

17x+y=85?

OpenStudy (anonymous):

17x + 10y = 85

OpenStudy (anonymous):

im sorry, i meant 17x+10y=85

OpenStudy (anonymous):

Yeah that's right :D

OpenStudy (anonymous):

Thank youuu